# Multidimensional constructs and moderated linear and nonlinear factor analysis

### v1

@@ Abstract
Multidimensional factor models with moderations on all model parameters have so far been limited to single-factor and two-factor models. This does not align well with existing psychological measures, which are commonly intended to assess 3-5 dimensions of a latent construct. In this paper, we introduce a penalized maximum likelihood approach for multidimensional MNLFA that permits moderation of item intercepts, loadings, residual variances, factor means, variances, and correlations across three or more latent factors. Our approach incorporates ridge, lasso, and alignment penalties to stabilize estimation and detect partial measurement non-invariance while preserving model interpretability. We derive closed-form analytic gradients of the likelihood, eliminating the need for costly numerical or MCMC-based approximations, and demonstrate how this dramatically improves computational efficiency. Through simulation and application, we illustrate that penalized MNLFA recovers complex moderation patterns in multidimensional constructs and provides a scalable alternative to Bayesian implementations. We conclude by discussing the theoretical implications of penalization for measurement invariance, computational considerations, and future directions for extending the framework to categorical indicators, longitudinal data, and applied research contexts.
@@ Introduction
The applications of moderated nonlinear factor analysis have been limited thus far to only two latent factors due to the inherent complexity of ensuring the latent variable covariance matrix is valid for all subgroups implied by the design matrix of moderating factors. The current approach for two-factor models is to induce on an unbounded continuum and then use Fisher’s z-transformation to restrict the parameter to the [-1,1] bound of correlations. Extending to more than two latent factors cannot utilize the same method using independent unbounded scores transformed to the [-1,1] scale because independent applications of Fisher’s z-transformations do not ensure the entire correlation matrix is valid. Any methods for inducing impact on individual correlations must simultaneously ensure each correlation is between [-1,1] and that the matrix of correlations is positive definite. Achieving these simultaneous constraints has been a longstanding problem in the simulation of correlation matrices (Joe,, 2006; Lewandowski et al., 2009a, ; Pourahmadi and Wang,, 2015; Archakov and Hansen,, 2021) and in covariance prediction (Pinheiro and Bates,, 1996; Pourahmadi,, 1999, 2011; Bucci et al.,, 2022). Modified conceptualizations of these methods are used in this article to address the dimensional limitation of moderated nonlinear factor analysis. We generalize moderated nonlinear factor analysis to flexibly allow moderation in the latent factor correlation parameters in models with three or more factors.
The remainder of this article is organized as follows. Next, we describe the MNLFA model in more detail for one- and two-dimensional factor models. Then, we described how MNLFA can be extended to $m$-dimensional factor models, which is the primary contribution of this article. We provide examples of how the general specification can induce a variety of correlational structures for specific subgroups. We then report on an application of the described multidimensional moderated nonlinear factor model for the study of psychopathology. We end with a discussion of the model and potential limitations and areas of future study.
@@
We focus this article on latent variable models with continuous indicators, but these methods generalize to binary and Likert-type responses with some modifications. To begin, consider the measurement model for items $i=1,\ldots,I$ with observed factor indicators $y$. The single factor latent variable $\eta$ for $n=1,\ldots,N$ cases can be described as
@@
The link function $g(.)$ depends on the scale of the observed indicators and for approximately continuous variables, we can simply use the identity link function. The major difference between the above model and a traditional single-factor model is the inclusion of moderation effects on model parameters (e.g, $\nu_{ni}=\nu_{0i}+\mathbf{x}_{n}^{t}\bm{\delta}_{\nu i}$ is the moderation on item intercepts). One way to conceptualize this approach to parameterizing the factor model is to think of hierarchical linear models where the intercepts and slopes (factor loadings) are outcomes of interest (see Raudenbush and Bryk,, 2002, p. 80). Shifting to a hierarchical/multi-level perspective for these models makes a direct connection between testing for non-invariance across a wide range of possible variables with well-known methods of linear modeling.
A major benefit of connecting evaluating non-invariance to linear modeling is the flexibility that allows for unique testing of potentially nonlinear relationships between person factors and non-invariance in parameters. For example, Molenaar, (2021) illustrated how generalized additive models could be used to evaluate non-invariance. This flexibility opens the door to evaluating even more complex interrelationships between model parameters and person characteristics through interactions among characteristics. Along with the benefits, evaluating non-invariance now shares many of the same concerns that arise with respect to collinearity in the design matrix for the predictors and overall model complexity. We will return to discuss collinearity and model complexity when we discuss priors later.
@@
Fisher’s z-transformation (Eq. 11) provides a straightforward method for inducing varying factor correlations separately from differences in factor variances. However, to extend the model to three or more dimensions in the latent factor structure, Fisher’s z-transformation approach does not work to jointly ensure the positive definiteness of the latent variable covariance matrix. Next, we introduce an approach to ensuring the positive definiteness that could be used in m-dimensional latent variable space.
@@
Extending beyond two latent factors effectively is a complex balancing act of where in the model do you allow moderation to be estimated. A fully saturated multidimensional model has moderation in all parameters–factor loadings, item intercepts, item residual variances, factor means, factor variances, and factor correlations. However, as we discuss in more detail in the Estimation & Empirical Identification section, the fully saturated model may not be estimable in all situations due to limitations of the data and model at hand. But first, we describe the fully saturated model and the process for inducing moderation in the factor correlation matrix to illustrate the potential for these models to explore a wide range of possible sources of measurement non-invariance in different parts of the measurement model. Moderation in the item parameters is relatively straightforward and can be shown as below
@@
The moderation in the item parameters is achieved through a model specification similar to that of the single or two-factor model. A new wrinkle involves specifying how moderation is defined in the factor loading matrix to account for possible cross-loadings. Simple structure can be assumed where each item loads onto only one factor, or any number of cross-loadings can be specified. The simplest solution is to fix all non-theorized cross-loadings to zero. A more general solution is to freely estimate a subset of loadings which we believe primarily load on different factors, and then use regularizing priors on cross-loadings. The complexity of extending to greater than two factors arises with the person-specific factor covariance matrices ($\bm{\Phi}_{n}$).
@@
Inducing impact, or moderation, in Cholesky factors can be accomplished using a partial correlation parameterization (Joe,, 2006; Lewandowski et al., 2009b, ) or a hypersphere parameterization (Pinheiro and Bates,, 1996; Rapisarda et al.,, 2007; Pourahmadi and Wang,, 2015). First, consider a vector of unconstrained parameters $\bm{\theta}_{n}$ with support of any real number. For $M$-factors, the number elements in $\bm{\gamma}_{n}$ is $M\times(M-1)/2=M^{\prime}$. Using the partial correlation parameterization, the vector $\bm{\gamma}_{n}$ with element $\gamma_{n12},\gamma_{n13},...,\gamma_{n1M},...,\gamma_{nM(M-1)}]$ can be used to construct the Cholesky factorization of the factor correlation matrix by first transforming to partial correlations using the hyperbolic transformation, $\theta^{\ast}=\text{tanh}\left(\gamma\right)=\frac{\exp(2\gamma)-1}{\exp(2\gamma)+1}$, resulting in the intermediary constraint of [1,1][-1,1] which partially leads to the interpretation as partial correlations. The Cholesky factorization is then built from this vector of partial correlations using the following two steps, and for the sake of simplicity, we drop the person subscript $n$.
@@
Estimating moderated nonlinear factor analyses is an active area of methodological research (Brandt et al.,, 2023, 2025; Chen and Bauer,, 2024; Gottfredson et al.,, 2019; Bauer et al.,, 2020; Kolbe et al.,, 2024). A common theme across these studies involved the use of regularization to help identify the model given the complexity of the likelihood. An approach with promise involves using Bayesian methods to approximate the posterior distribution and regularizing priors such as spike-and-slab priors for the DIF parameters (Brandt et al.,, 2023, 2025; Chen and Bauer,, 2024). The R package (Belzak,, 2023) implements Bayesian moderated nonlinear factor analyses for models with up to two latent factors relying on regularizing priors for DIF parameters (Belzak,, 2023) regDIF uses the general Bayesian MCMC package Just Another Gibbs Sampler (JAGS) to sample the marginal posterior distributions. JAGS is incredible powerful and flexible, but can be painfully slow for models with large numbers of parameters. To help overcome this limitation, we implemented the multidimensional moderated nonlinear factor analysis in Stan (REF), which is theoretically more efficient and can more easily explore the complex posterior distributions. After getting a working version of the full multidimensional MNLFA model in Stan, we ran into the same problem as with JAGS: it was painfully slow, still taking days to estimate the model with reasonable sample sizes.
@@
The severe computational time led to exploring alternative estimation approaches and identifying where the bottleneck in computation is in estimating the model. The reason the model in Stan was taking so long is that, under the hood, Stan is approximating the gradient, the first derivative of the posterior, for each model parameter, and then moving along this approximated curve to get to the next position in the posterior surface. This gradient approximation was very computationally demanding in a general multidimensional MNLFA model, leading to the excessive slowness we observed. To overcome this complexity, we sought an estimation approach that didn’t require a numerical approximation of the gradient. We returned to old school maximum likelihood methods. Because the MNLFA model assumes a multivariate normal distribution for the responses, the likelihood function is the well-known multivariate normal, and as it turns out, the gradients for all model parameters in the MNLFA have a closed form that we can make use of when estimating the model.
@@
Traditional methods of maximum likelihood do not incorporate regularization, but we can incorporate penalization for complexity that acts as a regularizing factor (Asparouhov and Muthén,, 2024). Instead of directly working with the log-likelihood, we can optimize a weighted composite of the log-likelihood ($L(.)$) and a penalty ($P(.)$),
$L_{w}(\theta)$ $=(1-w_{0})L(\theta)-w_{0}P(\theta),$ (21)
@@
where $w_{0}$ is the weight given to the "prior" or the penalty term, $\theta$ is the vector of all parameters being estimated, and $L_{w}(\theta)$ represents the composite log-likelihood. The penalty term can take different forms, but the method we propose is based on a splitting the parameter vector into pieces $(\theta_{0},\theta_{1})$ where $\theta_{0}$ are the non-regularized terms (e.g., baseline item intercepts, factor loadings, etc.) and where $\theta_{1}$ contains the DIF/impact parameters. The penalty is then placed only with respect to $\theta_{1}$. A few such penalties can be defined by summing up all the squared differences among the parameters in $\theta_{1}$,
@@
The penalized log-likelihood itself is relatively straightforward to set up to maximize using numerical methods, such as using routine optimization routines in any statistical package (e.g., optim(.) or nlminb(.) in R). However, MNLFA requires using a person-wise constructed mean and covariance matrix when evaluating the multivariate normal density function. This creates a computationally complex situation where we cannot use sufficient statistics to simplify the overall likelihood expression, as is done in lavaan. Instead must evaluate the density function for each case, or person, in our dataset after constructing the person-specific model implied mean and covariance matrix for the observed data. This complexity issue is similar to those found when trying to set up routines to estimate a multilevel structural equation model using full information maximum likelihood (see Rosseel,, 2021, for a deeper discussion on this topic). However, the tricks for efficiency described by Rosseel, (2021) may not be usable to improve the efficiency of evaluating the likelihood in MNLFA because no two persons in a dataset may have the exact same combination of moderating variables ($X$) or missing data pattern; though I certainly could be mistaken about this limitation. Due to this constraint, the only major source for efficiency gains in estimating the model lies with how the gradient of the log-likelihood is evaluated.
@@
Hessian approximation
The Hessian, or observed information matrix, of the parameter vector, $\bm{\theta}$, is a necessary contribution to approximate the uncertainty in the estimate parameters. We recommend using a sandwich-like estimator for the standard errors using the inverse of the observed information matrix and the outer-product of the gradient (Freedman,, 2006).

### v2

@@ Abstract
Multidimensional factor models with moderations on all model parameters have so far been limited to single-factor and two-factor models. This does not align well with existing psychological measures, which are commonly intended to assess 3-5 dimensions of a latent construct. In this paper, I introduce a penalized maximum likelihood approach for multidimensional MNLFA that permits moderation of item intercepts, loadings, residual variances, factor means, variances, and correlations across three or more latent factors. Our approach incorporates ridge, lasso, and alignment penalties to stabilize estimation and detect partial measurement non-invariance while preserving model interpretability. I derive closed-form analytic gradients of the likelihood, eliminating the need for costly numerical or MCMC-based approximations, and demonstrate how this dramatically improves computational efficiency. Through simulation and application, I illustrate that penalized MNLFA recovers complex moderation patterns in multidimensional constructs and provides a scalable alternative to Bayesian implementations. I conclude by discussing the theoretical implications of penalization for measurement invariance, computational considerations, and future directions for extending the framework to categorical indicators, longitudinal data, and applied research contexts.
@@ Introduction
The applications of moderated nonlinear factor analysis have been limited thus far to only two latent factors due to the inherent complexity of ensuring the latent variable covariance matrix is valid for all subgroups implied by the design matrix of moderating factors. The current approach for two-factor models is to induce on an unbounded continuum and then use Fisher’s z-transformation to restrict the parameter to the [-1,1] bound of correlations. Extending to more than two latent factors cannot utilize the same method using independent unbounded scores transformed to the [-1,1] scale because independent applications of Fisher’s z-transformations do not ensure the entire correlation matrix is valid. Any methods for inducing impact on individual correlations must simultaneously ensure each correlation is between [-1,1] and that the matrix of correlations is positive definite. Achieving these simultaneous constraints has been a longstanding problem in the simulation of correlation matrices (Joe,, 2006; Lewandowski et al., 2009a, ; Pourahmadi and Wang,, 2015; Archakov and Hansen,, 2021) and in covariance prediction (Pinheiro and Bates,, 1996; Pourahmadi,, 1999, 2011; Bucci et al.,, 2022). Modified conceptualizations of these methods are used in this article to address the dimensional limitation of moderated nonlinear factor analysis. I generalize moderated nonlinear factor analysis to flexibly allow moderation in the latent factor correlation parameters in models with three or more factors.
The remainder of this article is organized as follows. Next, I describe the MNLFA model in more detail for one- and two-dimensional factor models. Then, I described how MNLFA can be extended to $m$-dimensional factor models, which is the primary contribution of this article. I provide examples of how the general specification can induce a variety of correlational structures for specific subgroups. I then report on an application of the described multidimensional moderated nonlinear factor model for the study of psychopathology. I end with a discussion of the model and potential limitations and areas of future study.
@@
I focus this article on latent variable models with continuous indicators, but these methods generalize to binary and Likert-type responses with some modifications. To begin, consider the measurement model for items $i=1,\ldots,I$ with observed factor indicators $y$. The single factor latent variable $\eta$ for $n=1,\ldots,N$ cases can be described as
@@
The link function $g(.)$ depends on the scale of the observed indicators and for approximately continuous variables, I can simply use the identity link function. The major difference between the above model and a traditional single-factor model is the inclusion of moderation effects on model parameters (e.g, $\nu_{ni}=\nu_{0i}+\mathbf{x}_{n}^{t}\bm{\delta}_{\nu i}$ is the moderation on item intercepts). One way to conceptualize this approach to parameterizing the factor model is to think of hierarchical linear models where the intercepts and slopes (factor loadings) are outcomes of interest (see Raudenbush and Bryk,, 2002, p. 80). Shifting to a hierarchical/multi-level perspective for these models makes a direct connection between testing for non-invariance across a wide range of possible variables with well-known methods of linear modeling.
A major benefit of connecting evaluating non-invariance to linear modeling is the flexibility that allows for unique testing of potentially nonlinear relationships between person factors and non-invariance in parameters. For example, Molenaar, (2021) illustrated how generalized additive models could be used to evaluate non-invariance. This flexibility opens the door to evaluating even more complex interrelationships between model parameters and person characteristics through interactions among characteristics. Along with the benefits, evaluating non-invariance now shares many of the same concerns that arise with respect to collinearity in the design matrix for the predictors and overall model complexity. I will return to discuss collinearity and model complexity when I discuss priors later.
@@
Fisher’s z-transformation (Eq. 11) provides a straightforward method for inducing varying factor correlations separately from differences in factor variances. However, to extend the model to three or more dimensions in the latent factor structure, Fisher’s z-transformation approach does not work to jointly ensure the positive definiteness of the latent variable covariance matrix. Next, I introduce an approach to ensuring the positive definiteness that could be used in m-dimensional latent variable space.
@@
Extending beyond two latent factors effectively is a complex balancing act of where in the model do you allow moderation to be estimated. A fully saturated multidimensional model has moderation in all parameters–factor loadings, item intercepts, item residual variances, factor means, factor variances, and factor correlations. However, as I discuss in more detail in the Estimation & Empirical Identification section, the fully saturated model may not be estimable in all situations due to limitations of the data and model at hand. But first, I describe the fully saturated model and the process for inducing moderation in the factor correlation matrix to illustrate the potential for these models to explore a wide range of possible sources of measurement non-invariance in different parts of the measurement model. Moderation in the item parameters is relatively straightforward and can be shown as below
@@
The moderation in the item parameters is achieved through a model specification similar to that of the single or two-factor model. A new wrinkle involves specifying how moderation is defined in the factor loading matrix to account for possible cross-loadings. Simple structure can be assumed where each item loads onto only one factor, or any number of cross-loadings can be specified. The simplest solution is to fix all non-theorized cross-loadings to zero. A more general solution is to freely estimate a subset of loadings which I believe primarily load on different factors, and then use regularizing priors on cross-loadings. The complexity of extending to greater than two factors arises with the person-specific factor covariance matrices ($\bm{\Phi}_{n}$).
@@
Inducing impact, or moderation, in Cholesky factors can be accomplished using a partial correlation parameterization (Joe,, 2006; Lewandowski et al., 2009b, ) or a hypersphere parameterization (Pinheiro and Bates,, 1996; Rapisarda et al.,, 2007; Pourahmadi and Wang,, 2015). First, consider a vector of unconstrained parameters $\bm{\theta}_{n}$ with support of any real number. For $M$-factors, the number elements in $\bm{\gamma}_{n}$ is $M\times(M-1)/2=M^{\prime}$. Using the partial correlation parameterization, the vector $\bm{\gamma}_{n}$ with element $\gamma_{n12},\gamma_{n13},...,\gamma_{n1M},...,\gamma_{nM(M-1)}]$ can be used to construct the Cholesky factorization of the factor correlation matrix by first transforming to partial correlations using the hyperbolic transformation, $\theta^{\ast}=\text{tanh}\left(\gamma\right)=\frac{\exp(2\gamma)-1}{\exp(2\gamma)+1}$, resulting in the intermediary constraint of [1,1][-1,1] which partially leads to the interpretation as partial correlations. The Cholesky factorization is then built from this vector of partial correlations using the following two steps, and for the sake of simplicity, I drop the person subscript $n$.
@@
#’ R Code
#’ Builds a Cholesky factor (lower triangular matrix of dimension K) from a vector of unconstrained parameters (length K*(K-1)/2)
#’ @param x, vector of unconstrainted partial correlations length(x) must equal K*(K-1)/2
#’ @param K, dimension of matrix (KxK)
construct_chol_mat_lkj <- function(x,K){
# validation check
if(length(x) != K*(K-1)/2){
warning(paste0("Number of supplied unconstrained partial correlations ’x’ not valid for supplied ’K’. Length of x=", length(x),", but should equal K*(K-1)/2 = ",K*(K-1)/2))
z <- matrix(0, nrow=K,ncol=K) ;
w <- matrix(0, nrow=K,ncol=K) ;
r <- matrix(0, nrow=K,ncol=K) ;
x_tanh <- numeric(length(x));
x_tanh <- tanh(x); #hyperbolic tan tranform of y (unconstrained partial correlations)
z[lower.tri(z,FALSE)] <- x_tanh;
diag(z) = 1.0;
for(i in 1:K){
for(j in 1:K){
if(i < j) w[i,j] = 0.0;
if(1 == j && j == i) w[i,j] = 1.0;
if(1 == j && j < i) w[i,j] = z[i,j];
if(1 < j && j == i){
w[i,j] = cumul_prod( (1.0 - (z[i,1:(j-1)])^2.0)^0.5 );
if(1 < j && j < i){
w[i,j] = z[i,j] * cumul_prod( (1.0 - (z[i,1:(j-1)])^2.0)^0.5 );
w;
#’ Compute cumulative product of a vector x
#’ @param x a numeric vector (e.g., c(1,2,3,4))
#’ @returns a numeric value of the resulting cumulative product of x.
cumul_prod <- function(x){
K = length(x);
z = 1;
for(k in 1:K){
z = z*(x[k]);
return(z);
#’ Examples:
# 2x2 matrix => a single correlation
corr_mat <- matrix(c(1,0.70,0.70,1.0),ncol=2)
corr_mat
x <- atanh(0.70) # inverse Fisher transform correlation to get input for cholesky construction
K <- 2
L <- construct_chol_mat_lkj(x, K)
L %*% t(L)
all.equal(L %*% t(L) , corr_mat)
@@
Estimating moderated nonlinear factor analyses is an active area of methodological research (Brandt et al.,, 2023, 2025; Chen and Bauer,, 2024; Gottfredson et al.,, 2019; Bauer et al.,, 2020; Kolbe et al.,, 2024). A common theme across these studies involved the use of regularization to help identify the model given the complexity of the likelihood. An approach with promise involves using Bayesian methods to approximate the posterior distribution and regularizing priors such as spike-and-slab priors for the DIF parameters (Brandt et al.,, 2023, 2025; Chen and Bauer,, 2024). The R package (Belzak,, 2023) implements Bayesian moderated nonlinear factor analyses for models with up to two latent factors relying on regularizing priors for DIF parameters (Belzak,, 2023) regDIF uses the general Bayesian MCMC package Just Another Gibbs Sampler (JAGS) to sample the marginal posterior distributions. JAGS is incredible powerful and flexible, but can be painfully slow for models with large numbers of parameters. To help overcome this limitation, I implemented the multidimensional moderated nonlinear factor analysis in Stan (Carpenter et al.,, 2017), which is theoretically more efficient and can more easily explore the complex posterior distributions.
To help overcome this limitation, I implemented the multidimensional moderated nonlinear factor analysis in Stan (Carpenter et al.,, 2017), which is theoretically more efficient and can more easily explore the complex posterior distributions.
I implemented the multidimensional MNLFA model in Stan (described next) and then, as I discuss in the Bayesian estimation section below, ended up trying to also implement a penalized maximum likelihood approach due to computational issues with the first Bayesian method implemented.
Bayesian approach
A Bayesian latent variable model can be implemented in several different ways depending on how the likelihood is set-up. Initially, I built the multidimensional MNLFA using a conditional likelihood approach to bring the implementation in line with how the package ’brms’ is likely to implement latent variable models in version 3 (https://github.com/paul-buerkner/brms/issues/304).
The likelihood is implemented as layed out in Figure 1. A complexity when implementing the model in Stan is the need to create independent foundation parameters that are then transformed into the parameters necessary for the likelihood function contribution itself. Also known as a non-centered parameterization so that the parameters being directly sampled have no direct dependence. This is accomplished largely by using independent normal parameters that are then transformed to get the quantities of interest. For instance, assuming there are $N$ respondents and $M$ latent variables, then the non-centered parameterization for the latent factors can be implemented as:
# Stan code
data {
int N; //number of cases
int P; //number of observed variables
int M; //number of latent factors
array[N] vector[P] y_obs;
parameters {
array[N*M] real eta_raw; //independent latent factors (MVN)
... //other parameters
transformed parametrs {
array[N] vector[M] eta; //rescaled factor scores
for(n in 1:N){
//Alpha = moderated factor means (Mx1)
//Psi_chol, moderated factor correlation matrix-cholesky parameterization (lower MxM)
//Psi_var, moderated factor variances (Mx1)
//Lambda, moderated factor loading matrix (PxM)
//Nu, moderated item intercepts (Px1)
//Theta, diagonal matrix of moderated item residual variances (PxP)
.//code to get moderated parameters omitted for simplificity
eta[n] = Alpha + to_vector((Psi_chol * to_vector(eta_raw[((n-1)*M+1):(n*M)]) ) * diag_matrix(Psi_var)); // non-centered parameterization of latent factor scores
implied_mean[n] = Nu + Lambda*eta;
implied_residual_cov[n] = Theta;
model {
for(n in 1:N){
//to help conserve memory, the transformed parameters could be constructed within this loop instead of the transformed parameters block
target += multi_normal_chol_lpdf(y[n] | implied_mean[n], implied_residual_cov[n]);
target += std_normal_lpdf(eta_raw);
.//other priors
This approach becomes even more memory intensive for item factor analysis (MNLFA with binary/ordinal items) due to the need to supply ‘nuisance” parameters $u$ to utilize the underlying latent response distribution of the probit model.
Marginal likelihood approach
However, this led to several issues when trying to implement it. First, it was PAINFULLY slow, partially due to the huge number of parameters (number of cases $\times$ number of latent factors). This becomes worse in the case of binary items (probit model) where a nusance parameter is used under the hood for each response (not just each person). To help overcome this limitation, I rewrote the model to utilize the marginal likelihood, similar to what is utilized in lavaan (Rosseel,, 2012) for estimating simple factor models. This led to something that was seemingly useful. Secondly, the conditional likelihood approach led to near impossible computer memory constraints. When trying the implement the model above, I could not get the model to run with sample sizes above 5000 (individual cases, not individual responses) without requesting over 1000 GB of ram on a cluster computer.
Although what follows builds on the Bayesian implementation described above, I did not do this until after trying the penalized maximum likelihood approach described in the next major section.
The marginal likelihood implemented in Stan is as follows. First, consider the observed factor indicators as continuous variables, then person-wise marginal likelihood can be expressed as using the notation below.
$\bm{y}_{n}$ $\sim\text{MVN}\left(\bm{\nu}_{n}+\bm{\Lambda}_{n}\bm{\alpha}_{n},\bm{\Lambda}_{n}\bm{\Phi}_{n}\bm{\Lambda}_{n}^{t}+\bm{\Theta}_{n}\right)$ (21)
$\nu_{n,i}$ $=\nu_{0i}+\mathbf{x}_{n}^{t}\bm{\delta}_{\nu i}$ (22)
$\lambda_{n,im}$ $=\lambda_{0im}+\mathbf{x}_{n}^{t}\bm{\delta}_{\lambda im}$ (23)
$\theta_{0,i}$ $=\theta_{0i}\exp(\mathbf{x}_{n}^{t}\bm{\delta}_{\theta i})$ (24)
$\alpha_{n,m}$ $=\alpha_{0m}+\mathbf{x}_{n}^{t}\bm{\delta}_{\alpha m}$ (25)
$\bm{\Phi}_{n}$ $=\left(\bm{\Phi}_{n}^{L}\right)^{t}\left(\bm{\Phi}_{n}^{D}\right)\left(\bm{\Phi}_{n}^{L}\right)$ (26)
$\bm{\Phi}_{n}^{D}$ $=\text{diag}(\phi_{n1},...,\phi_{nm})$ (27)
$\phi_{n,m}$ $=\phi_{0m}\exp(\mathbf{x}_{n}^{t}\bm{\delta}_{\phi m})$ (28)
$\bm{\Phi}_{n}^{L}$ $=f(\bm{\gamma}_{n})$ (29)
$\gamma_{n,m^{\prime}}$ $=\gamma_{nm^{\prime}}+\mathbf{x}_{n}^{t}\bm{\delta}_{\gamma m^{\prime}}$ (30)
Using this framing of the likelihood, the dimension of the parameter space does not increase with sample size leading to some efficiency gains. A major benefit of this approach is that common SEM/CFA rules for identification (Bollen,, 1989) can be used to help ensure the baseline parameters are easily identified which then help to empirically identify the remaining parameters.
For example, constraining the baseline factor loadings of each variable/item’s primary factor loading to be positive. This can be accomplished using a constraint on the parameter itself, or using a semi-informative prior that does not put much, if any, weight below zero; e.g., $\lambda_{0,im}\sim\text{Normal}(1.0,0.5)$.
The factor means/intercepts can be fixed to 0.0, then allow for impact through non-zero coefficients in $\bm{\delta}_{\alpha}$.
A fixed baseline factor variance is also possible. Such as fixing the factor variance to 1.0, and allowing the factor variance to vary as a function of background characteristics. This can work well in conjunction with a semi-informative prior on the factor loadings to establish the scale of the factor variables while still allowing for population differences in scale as a function of background characteristics. These differences would then be interpretable as changes in scales compared to the average across populations.
Spike-and-slab priors for DIF & Impact
Spike-and-slab priors have seen a growing interest in the area of DIF detection (Brandt et al.,, 2018; Belzak and Bauer,, 2020; Chen et al.,, 2021; Brandt et al.,, 2025). The spike-and-slab prior can be implemented in a variety of ways depending on the needs of the situation. One way previously implemented is
$\delta$ $\sim\phi\text{Normal}\left(0,\psi\tau^{2}\sigma^{2}\right)$ (31)
$\psi$ $\sim\text{exponential}\left(0.5\right)$ (32)
$\phi$ $\sim\text{Beta}\left(1,1\right)$ (33)
$\tau^{-1}$ $\sim\text{Gamma}\left(9,4\right)$ (34)
where $\delta$ is a placeholder for any of the DIF/impact effects, $\phi$ is the inclusion probability, $\psi$ and $\tau$ are regularizing variance parameters across a group of jointly regularized parameters, and $\sigma$ is the item specific residual variance.
When implementing the above parameterization of the spike-and-slab prior, I found that the individual regularizing variance parameters were not unique leading to sampling difficulties in Stan. An alternative parameterization designed specifically for Stan is described next. Thomson et al., (2019) provided a logit-normal continuous parameterization that is especially efficient for use in Stan since all sampled parameters are continuous with unbounded support that can then be transformed to the standard spike-and-slab structure.
$\delta$ $\sim\phi\tau\delta_{0}$ (35)
$\delta_{0}$ $\sim\text{Normal}\left(0,1\right)$ (36)
$\phi$ $=\text{inv-logit}\left(\phi_{0}\right)$ (37)
$\phi_{0}$ $=\log\left(\frac{\phi_{\mu}}{1-\phi_{\mu}}\right)+\text{Normal}\left(0,1\right)\times\phi_{sd}$ (38)
$\tau$ $=\tau_{0}$ (39)
While the above prior parameterization has $\tau$ fixed, as done in all the simulations conducted by Thompson and colleagues which greatly improves speed and sampling effeciency of the spike and slab prior. The variance can be given a hyper-prior, such as a Gamma distribution, to provide an “adaptive” component. However, in testing, I’ve found that allowing the the “adaptive” piece simply leads to significant increase in computation time and more diffuse posterior distributions for the DIF/impact effects. The major advantage of this approach is only needing to use standard normal priors then shifting/rescaling as needed to get the inclusion probability shifted to be more or less regularizing (using prior inclusion probability mean, $\phi_{\mu}$, and standard deviation, $\phi_{sd}$), and increasing or decreasing $\tau_{0}$ for more or less regularizing. Together, this leads to a more interpretable regularizing prior for me that is easy to specify and modify as needed for sensitivity analyses.
# Stan code
data {
int Q;//number of regularied parameters
real pen_mu ; // pre-transformed IP mean on logit scale
real<lower=0> pen_sd ;
real<lower=0> pen_tau;//if a fixed value
parameters {
array[Q] real delta_raw;
real pen_phi;
//array[Q] real<lower=0> pen_tau;
transformed parameters {
real delta;
real pen_inclusion_prob;
pen_inclusion_prob = inv_logit(pen_mu + pen_sd .* to_vector(pen_phi));
delta = detla_raw .* pen_tau .* pen_inclusion_prob;
model {
target += std_normal_lpdf(delta_raw); //dif/impact raw parameter
target += std_normal_lpdf(pen_phi); //inclusion probability
//target += gamma_lpdf(pen_tau | tau_a, tau_b); //only used if allowing for varying penalty variance
Penalized maximum likelihood
@@
The severe computational time led in the conditional likelihood approach described in the Bayesian section above, I decided to explore alternative estimation approaches and identifying where the bottleneck in computation is in estimating the model. The reason the model in Stan was taking so long is that, under the hood, Stan is approximating the gradient, the first derivative of the posterior, for each model parameter, and then moving along this approximated curve to get to the next position in the posterior surface. This gradient approximation was very computationally demanding in a general multidimensional MNLFA model, leading to the excessive slowness I observed. To overcome this complexity, I sought an estimation approach that didn’t require a numerical approximation of the gradient. I returned to old school maximum likelihood methods. Because the MNLFA model assumes a multivariate normal distribution for the responses, the likelihood function is the well-known multivariate normal, and as it turns out, the gradients for all model parameters in the MNLFA have a closed form that I can make use of when estimating the model.
@@
Traditional methods of maximum likelihood do not incorporate regularization, but I can incorporate penalization for complexity that acts as a regularizing factor (Asparouhov and Muthén,, 2024). Instead of directly working with the log-likelihood, I can optimize a weighted composite of the log-likelihood ($L(.)$) and a penalty ($P(.)$),
$L_{w}(\theta)$ $=L(\theta)-w_{0}P(\theta),$ (40)
@@
where $w_{0}$ is the weight given to the "prior" or the penalty term, $\theta$ is the vector of all parameters being estimated, and $L_{w}(\theta)$ represents the composite log-likelihood. The penalty term can take different forms, but the method I propose is based on a splitting the parameter vector into pieces $(\theta_{0},\theta_{1})$ where $\theta_{0}$ are the non-regularized terms (e.g., baseline item intercepts, factor loadings, etc.) and where $\theta_{1}$ contains the DIF/impact parameters. The penalty is then placed only with respect to $\theta_{1}$. A few such penalties can be defined by summing up all the squared differences among the parameters in $\theta_{1}$,
@@
The penalized log-likelihood itself is relatively straightforward to set up to maximize using numerical methods, such as using routine optimization routines in any statistical package (e.g., optim(.) or nlminb(.) in R). However, MNLFA requires using a person-wise constructed mean and covariance matrix when evaluating the multivariate normal density function. This creates a computationally complex situation where I cannot use sufficient statistics to simplify the overall likelihood expression, as is done in lavaan. Instead must evaluate the density function for each case, or person, in our dataset after constructing the person-specific model implied mean and covariance matrix for the observed data. This complexity issue is similar to those found when trying to set up routines to estimate a multilevel structural equation model using full information maximum likelihood (see Rosseel,, 2021, for a deeper discussion on this topic). However, the tricks for efficiency described by Rosseel, (2021) may not be able to improve the efficiency of evaluating the likelihood in MNLFA because no two persons in a dataset may have the exact same combination of moderating variables ($X$) or missing data pattern; though I certainly could be mistaken about this limitation and future work should seek explore such possibilities. Due to this constraint, the only major source for efficiency gains in estimating the model lies with how the gradient of the log-likelihood is evaluated.
@@
Standard errors and the Hessian approximation
I recommend using a sandwich-like estimator for the standard errors using the inverse of the observed information matrix (Hessian) and the outer-product of the gradient (Freedman,, 2006). Obtaining an analytical definition for the Hessian is not straightforward, while approximating the Hessian has many existing (but slow) implementations. The Hessian, or observed information matrix, of the parameter vector, $\bm{\theta}$, is a necessary contribution to approximate the uncertainty in the estimated parameters. However, approximating this matrix is not straightforward, and though computationally intensive, the most straightforward approach is to use build in tools to approximate the hessian using numerical methods building on the analytical gradients described above.